OpenStudy (anonymous):

What are the approximate solutions of 3x2 − 5x = 3 rounded to the nearest hundredth?

OpenStudy (lgbasallote):

change to standard form first 3x^2 - 5x - 3 = 0 use quadratic formula to solve for x

OpenStudy (anonymous):

okay

OpenStudy (anikate):

or the x method which is also less wrting

OpenStudy (anonymous):

what is the x method?

OpenStudy (anikate):

let me show you

OpenStudy (anonymous):

okay

OpenStudy (anikate):

nvm just do quadratic formula for this, the x method cannot be used on this type of quadratic equation

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3x² - 5x - 3 = 0 Roots: 2.135041612, -0.4683749459 Discriminant: 61 Vertex: (0.8333333333, -5.083333333) Factored: f(x) = 3(x - 2.135041612)(x + 0.4683749459) Sum of Roots (-b/a): 1.666666666 Product of Roots (c/a): -1

OpenStudy (anonymous):

so would the roots be the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okaay